OpenStudy (anonymous):

Determine the equation of the parabola in the form y= a(x-p)^2 +pq a) the parabola has a vertex of (0,-3) and a x-intercept of 4 b) the x-intercept are x = -4 and x=2 and the graph is congruent to y= -2/3(x+1)^2 -3

OpenStudy (mathmate):

Please check for typo in the original equation: is it y= a(x-p)^2 +pq or y= a(x-p)^2 +q

OpenStudy (anonymous):

y=a(x-p)^2+q

OpenStudy (mathmate):

Since the vertex form of the parabola is y=a(x-h)^2+k where (h,k) is the vertex, we conclude that for y=a(x-p)^2+q with vertex at (0,-3), then p=0, q=-3, so y=ax^2-3 Since one zero is at x=4, and the zeroes of a parabola are symmetrical about the vertex, we conclude that there are two zeroes, x=4, and x=-4, resulting in the equation y=a(x-4)(x+4) = a(0)^2-3 Solving for a, a=-3/(-16)=3/16 => y=(3/16)x^2-3 b. y=-2/3(x+1)^2 -3 is already in the vertex form, a=-2/3 p=-1 q=-3

OpenStudy (anonymous):

thank u sooo much! :D

OpenStudy (mathmate):

You're welcome! :)